OpenStudy (anonymous):

Solve. Please show all of your work. The sum of twice a number and 4 less than the number is the same as the difference between -32 and the number. What is the number?

Parth (parthkohli):

Let the number be x: 2x - 4 = x + 32 x =32 + 4 x = 36

Parth (parthkohli):

Note that x + 32 is because x -(-32) = x + 32

OpenStudy (anonymous):

let the number be X 2X+ ( X-4)=X+32 3X-4=X+32 2X=36 X=18